OpenStudy (anonymous):

A physician saw 18 cases of chicken pox in the first week of an outbreak. Each week, he saw 6 more cases than the week before. After 10 weeks, how many cases of chicken pox did the physician see altogether? a. Write a sequence that shows the number of cases seen for each of the 10 weeks during the outbreak. b. Use Gauss’ method to find the sum. Be sure to show and explain your work.

OpenStudy (anonymous):

i need help with this answer

OpenStudy (anonymous):

If you can tell me what Gauss' method is...

OpenStudy (anonymous):

well i learned that guess methond is like for an example 1 to 100 something like that

OpenStudy (anonymous):

Gauss method is just a simple way of calculating values from a to b without making the mistake of writing out each number

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

ok well can someone explain me how to do this question ?

OpenStudy (anonymous):

a)18+10*6=18+60=78 cafter 10 weeks

OpenStudy (anonymous):

18 cases/week times 10 weeks = 180 cases

OpenStudy (anonymous):

ne additional group of six cases per week for 9 weeks.

OpenStudy (anonymous):

*one

OpenStudy (anonymous):

(n^2+n)/2 6((9^2+9)/2) 6((81+9)/2 3(90) 270 additional cases 270 + 180 = 450 cases total

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

What do you mean how? @Fazeelayaz

OpenStudy (anonymous):

why r u dividing it by 2

OpenStudy (anonymous):

Thats the formual

OpenStudy (anonymous):

*formula

OpenStudy (anonymous):

i agree with @zordoloom